OpenStudy (anonymous):

Evaluate the integral from -infinty to 0 (1/(2x-1)^3) and ttell whether it diverges or converges.

OpenStudy (anonymous):

\[\int\limits_{-\infty}^{0} 1/(2x-1)^{3}\]

OpenStudy (amistre64):

looks to be missing a 2

OpenStudy (amistre64):

\[\frac{-1}{2(2x-1)^2}\] is the integral part i believe

OpenStudy (amistre64):

\[\frac{-1}{2(2.0-1)^2}-(\frac{-1}{2(2.inf-1)^2})\] \[\frac{-1}{2(-1)^2}-(\frac{-1}{2(inf)^2})\] \[\frac{-1}{2}-(\frac{-1}{inf})\] \[\frac{-1}{2}-(0)\]

OpenStudy (across):

\[\int_{-\infty}^{0}\frac{1}{(2x-1)^3}dx=\left [ -\frac{1}{4(1-2x)^2} \right ]_{-\infty}^{0}\]\[\implies -\frac{1}{4}+0=-\frac{1}{4}.\]Seems to converge.

OpenStudy (amistre64):

4? really?

OpenStudy (across):

I performed the substitution\[u=2x-1,\]\[du=2dx.\]

OpenStudy (amistre64):

\[dy/dx\ \frac{-1}{2}(2x-1)^{-2} = \frac{2}{2}(2x-1)^{-3}\]

OpenStudy (amistre64):

opps, forgot the 2 that pops out of the inners :)

OpenStudy (amistre64):

so yeah, -1/4 is it ;)

OpenStudy (anonymous):

Thank You! :)